OpenStudy (ambermarie151):

The function c(f) = (f – 32) allows you to convert degrees Fahrenheit to degrees Celsius. Find the inverse of the function so that you can convert degrees Celsius back to degrees Fahrenheit. A. f(c) = c – 32 B. f(c) = c + 32 C. f(c) = c – 32 D. f(c) = c + 32

OpenStudy (ambermarie151):

@trojanpoem

OpenStudy (trojanpoem):

c(f) + 32 = f f = c(f) + 32 f(c) = c + 32